### v1

# Baillon-Bruck-Reich revisited: divergent-series parameters and strong convergence in the linear case

Abstract
The Krasnoselskii-Mann iteration is an important algorithm in optimization and variational analysis for finding fixed points of nonexpansive mappings. In the general case, it produces a sequence converging weakly to a fixed point provided the parameter sequence satisfies a divergent-series condition.
In this paper, we show that strong convergence holds provided the underlying nonexpansive mapping is linear. This improves on a celebrated result by Baillon, Bruck, and Reich from 1978, where the parameter sequence was assumed to be constant as well as on recent work where the parameters were bounded away from 0 and 11.
To Simeon Reich, on the occasion of his 80th birthday,
with deep thanks for decades of inspiration and mentorship.
2020 Mathematics Subject Classification: Primary 47H05, 47H09; Secondary 47N10, 65K05, 90C25.
Keywords: Baillon-Bruck-Reich theorem, fixed point, Hilbert space, Krasnoselskii-Mann iteration, nonexpansive mapping.
1 Introduction
Throughout this note,
$X$ is a real Hilbert space, (1)
with inner product $\left\langle{\cdot},{\cdot}\right\rangle$ and induced norm $\|\cdot\|$. We also assume that
$T\colon X\to X$ is nonexpansive, with $F:=\operatorname{Fix}T\neq\varnothing$. (2)
When $\lambda\in\mathbb{R}$, we set
$T_{\lambda}:=(1-\lambda)\operatorname{Id}+\lambda T,$ (3)
where $\operatorname{Id}\colon X\to X\colon x\mapsto x$ is the identity operator.
Many problems in optimization and variational analysis can be reduced to finding a point in $F$ (see, e.g., [2], [5], [9], and the references therein). A popular algorithmic approach is to employ the Krasnoselskii-Mann iteration for which we now state a basic convergence result:
Fact 1.1 (Reich).
Let $(\lambda_{n})_{n\in{\mathbb{N}}}$ be in $[0,1]$ and assume that $\sum_{n\in{\mathbb{N}}}(1-\lambda_{n})\lambda_{n}=+\infty$. Let $x_{0}\in X$. Then the sequence generated by
$(\forall{n\in{\mathbb{N}}})\quad x_{n+1}:=T_{\lambda_{n}}x_{n}$ (4)
converges weakly to a point in $F$. Moreover,
$x_{n}-Tx_{n}\to 0,$ (5)
and $(x_{n})_{n\in{\mathbb{N}}}$ is Fejér monotone with respect to $F$.
[…]
We refer the reader also to the recent book [6], dedicated entirely to Krasnoselskii-Mann iterations. When $T$ is linear, the Krasnoselskii-Mann iteration produces a strongly convergent sequence provided that the parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ is constant:
Fact 1.2 (Baillon-Bruck-Reich).
Suppose that $T$ is linear and that $\lambda\in\left]0,1\right[$. Let $x_{0}\in X$, and generate sequence $(x_{n})_{n\in{\mathbb{N}}}$ in $X$ via
$(\forall{n\in{\mathbb{N}}})\quad x_{n+1}:=T_{\lambda}x_{n}$ (6)
Then
$x_{n}\to P_{F}x_{0}.$ (7)
[…]
Recently, Section 1 was generalized in [3] to deal with a parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ satisfying $0<\inf_{n\in{\mathbb{N}}}\lambda_{n}\leq\sup_{n\in{\mathbb{N}}}\lambda_{n}<1$.
The goal of this paper is to extend the Baillon-Bruck-Reich result (Section 1) to the general parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ satisfying the divergent-series condition $\sum_{n\in{\mathbb{N}}}(1-\lambda_{n})\lambda_{n}=\infty$.
The remainder of the paper is organized as follows. In Section 2, we present the main result (Theorem 2.2) and its proof. The paper is concluded with various remarks.
The notation we employ is fairly standard and follows largely [2].
2 Main result
Proposition 2.1.
Recall that $T$ satisfies 2. Assume in addition that $T$ is linear. Then
$F^{\perp}=\overline{\operatorname{ran}}\,(\operatorname{Id}-T).$ (8)
Proof. Because $T$ is nonexpansive, we deduce from [2, Example 20.29] that $\operatorname{Id}-T$ is maximally monotone. In turn, it follows from [2, Example 20.17] that
$F=\ker(\operatorname{Id}-T)=\ker(\operatorname{Id}-T^{*}).$ (9)
On the other hand, because $\operatorname{Id}-T$ is a continuous linear operator, we have
$\big(\ker(\operatorname{Id}-T^{*})\big)^{\perp}=\overline{\operatorname{ran}}\,(\operatorname{Id}-T).$ (10)
[…]
We are now ready for the main result.
Theorem 2.2 (main result).
Recall that $T$ satisfies 2, and assume in addition that $T$ is linear. Let $(\lambda_{n})_{n\in{\mathbb{N}}}$ be in $[0,1]$ such that $\sum_{n\in{\mathbb{N}}}(1-\lambda_{n})\lambda_{n}=\infty$, and let $x_{0}\in X$. Then the sequence generated by
$(\forall{n\in{\mathbb{N}}})\quad x_{n+1}:=T_{\lambda_{n}}x_{n}$ (11)
converges strongly to $P_{F}x_{0}$.
Proof. Note that $F\subseteq\operatorname{Fix}T_{\lambda_{k}}$. Set $f_{0}:=P_{F}x_{0}$ and $g_{0}:=P_{F^{\perp}}x_{0}$ so that
$x_{0}=f_{0}+g_{0}.$ (12)
Clearly,
$x_{n+1}$ $=T_{\lambda_{n}}\cdots T_{\lambda_{0}}x_{0}=T_{\lambda_{n}}\cdots T_{\lambda_{0}}f_{0}+T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{0}=f_{0}+T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{0}=P_{F}x_{0}+T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{0}.$ (13)
It thus suffices to show that
$T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{0}\to 0.$ (14)
Because $0\in F\subseteq\operatorname{Fix}T_{\lambda_{k}}$, we have that $(\|T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{0}\|)_{n\in{\mathbb{N}}}$ is decreasing and so
$\ell:=\lim_{n\to\infty}\|T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{0}\|=\inf_{{n\in{\mathbb{N}}}}\|T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{0}\|$ (15)
exists. Next, 8 implies that $g_{0}\in F^{\perp}=\overline{\operatorname{ran}}\,(\operatorname{Id}-T)$. Let $\varepsilon>0$ and get $v_{\varepsilon}\in X$ such that $g_{\varepsilon}:=(\operatorname{Id}-T)v_{\varepsilon}\in{\operatorname{ran}}\,(\operatorname{Id}-T)$ and $\|g_{0}-g_{\varepsilon}\|\leq\varepsilon$. Now 5 from Section 1 yields
$T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{\varepsilon}=T_{\lambda_{n}}\cdots T_{\lambda_{0}}(\operatorname{Id}-T)v_{\varepsilon}=T_{\lambda_{n}}\cdots T_{\lambda_{0}}v_{\varepsilon}-TT_{\lambda_{n}}\cdots T_{\lambda_{0}}v_{\varepsilon}\to 0.$ (16)
Thus
$\ell$ $\leq\|T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{0}\|$ (by 15)
$\leq\|T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{0}-T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{\varepsilon}\|+\|T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{\varepsilon}\|$ (triangle inequality)
$\leq\|g_{0}-g_{\varepsilon}\|+\|T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{\varepsilon}\|$ (each $T_{\lambda_{k}}$ is nonexpansive)
$\leq\varepsilon+\|T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{\varepsilon}\|$ (choice of $v_{\varepsilon}$ and $g_{\varepsilon}$)
$\to\varepsilon\;\;\text{ as $n\to\infty$.}$ (by 16)
[…]
We conclude with some remarks concering Theorem $2.$
Remark 2.3 (without linearity, strong convergence fails).
Thanks to an example constructed by Genel and Lindenstrauss (see [7]), it is known that the sequence $(x_{n})_{n\in{\mathbb{N}}}$ generated by the Krasnoselskii-Mann iteration may fail to converge strongly even when $\lambda_{n}\equiv\tfrac{1}{2}$. A posteriori, the underlying nonexpansive mapping cannot be linear.
Remark 2.4 (without divergent-series parameters, convergence to a fixed point fails).
Suppose that $x_{0}\in X\smallsetminus\{0\}$, and consider the linear nonexpansive isometry $T=-\operatorname{Id}$. Then $F=\{0\}$. Assume that the parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ lies in $\left[0,\tfrac{1}{2}\right[$. Then the sequence $(x_{n})_{n\in{\mathbb{N}}}$ generated by the Krasnoselskii-Mann iteration satisfies
$x_{n}=\Big(\prod_{k=0}^{n-1}(1-2\lambda_{k})\Big)x_{0};$ (17)
moreover, $x_{n}\to 0$ $\Leftrightarrow$ $\sum_{n\in{\mathbb{N}}}\lambda_{k}=\infty$. Therefore, if $\sum_{n\in{\mathbb{N}}}(1-\lambda_{n})\lambda_{n}<\infty$, then $(x_{n})_{n\in{\mathbb{N}}}$ does not converge to a point in $F$.
Remark 2.5 (extension to the affine case).
If $T\colon X\to X$ is assumed to be affine, still with $F=\operatorname{Fix}T\neq\varnothing$, then Theorem 2.2 holds as well. The (standard translation) argument is outlined in [3, Section 3.2].
Remark 2.6 (for averaged $T$, one may increase the range of the parameter sequence).
If it is a priori known that $T$ is averaged, then one may enlarge the range of parameter sequence to go beyond 11. (See [3, Section 3.1] for how this is done.)
Remark 2.7 (an open problem).
The original paper by Baillon, Bruck, and Reich [1] was situated in a more general setting with respect to the underlying space and the nonexpansive map; in fact, $T$ was merely assumed to be odd rather than linear. At present, we do not know if it is possible to extend Theorem 2.2 from linear to odd nonexpansive mappings.
Acknowledgment
The research of SB was partially supported by a collaboration grant for mathematicians of the Simons Foundation. The research of HHB was partially supported by a Discovery Grant of the Natural Sciences and Engineering Research Council of Canada.

### v2

@@
 Abstract
-The Krasnoselskii-Mann iteration is an important algorithm in optimization and variational analysis for finding fixed points of nonexpansive mappings.
+The Krasnosel’skiĭ-Mann iteration is an important algorithm in optimization and variational analysis for finding fixed points of nonexpansive mappings.
 In the general case, it produces a sequence converging weakly to a fixed point provided the parameter sequence satisfies a divergent-series condition.
@@
 2020 Mathematics Subject Classification: Primary 47H05, 47H09; Secondary 47N10, 65K05, 90C25.
-Keywords: Baillon-Bruck-Reich theorem, fixed point, Hilbert space, Krasnoselskii-Mann iteration, nonexpansive mapping.
+Keywords: Baillon-Bruck-Reich theorem, fixed point, Hilbert space, Krasnosel’skiĭ-Mann iteration, nonexpansive mapping.
 1 Introduction
@@
 Many problems in optimization and variational analysis can be reduced to finding a point in $F$ (see, e.g., [2], [6], [10], and the references therein).
-A popular algorithmic approach is to employ the Krasnoselskii-Mann iteration for which we now state a basic convergence result:
+A popular algorithmic approach is to employ the Krasnosel’skiĭ-Mann iteration for which we now state a basic convergence result:
 Fact 1.1 (Reich).
@@
 […]
-We refer the reader also to the recent book [6], dedicated entirely to Krasnoselskii-Mann iterations.
-When $T$ is linear, the Krasnoselskii-Mann iteration produces a strongly convergent sequence provided that the parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ is constant:
+We refer the reader also to the recent book [7], dedicated entirely to Krasnosel’skiĭ-Mann iterations.
+When $T$ is linear, the Krasnosel’skiĭ-Mann iteration produces a strongly convergent sequence provided that the parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ is constant:
 Fact 1.2 (Baillon-Bruck-Reich).
@@
 The goal of this paper is to extend the Baillon-Bruck-Reich result (Section 1) to the general parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ satisfying the divergent-series condition $\sum_{n\in{\mathbb{N}}}(1-\lambda_{n})\lambda_{n}=\infty$.
-The remainder of the paper is organized as follows. In Section 2, we present the main result (Theorem 2.2) and its proof. The paper is concluded with various remarks.
+The remainder of the paper is organized as follows. In Section 2, we present the main result (Theorem 2.3), its proof, and corresponding comments. In Section 3, we offer some concluding remarks.
 The notation we employ is fairly standard and follows largely [2].
@@
 […]
+Remark $2.$
+We are grateful to a reviewer who outlined the following alternative proof of Section 2: Let $A\colon X\rightrightarrows X$ be a maximally monotone linear relation.
+Combining [4, Fact 2.1.(v) and Theorem 3.2] yields $(\ker A)^{\perp}=\overline{\operatorname{ran}}\,A^{*}=\overline{\operatorname{ran}}\,A$.
+Then Section 2 follows when one sets $A:=\operatorname{Id}-T$.
 We are now ready for the main result.
@@
 Now 5 from Section 1 yields
-$T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{\varepsilon}=T_{\lambda_{n}}\cdots T_{\lambda_{0}}(\operatorname{Id}-T)v_{\varepsilon}=T_{\lambda_{n}}\cdots T_{\lambda_{0}}v_{\varepsilon}-TT_{\lambda_{n}}\cdots T_{\lambda_{0}}v_{\varepsilon}\to 0.$ (16)
-Thus
+$T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{\varepsilon}=T_{\lambda_{n}}\cdots T_{\lambda_{0}}(\operatorname{Id}-T)v_{\varepsilon}=T_{\lambda_{n}}\cdots T_{\lambda_{0}}v_{\varepsilon}-TT_{\lambda_{n}}\cdots T_{\lambda_{0}}v_{\varepsilon}\to 0,$ (16)
+where we also used the fact that $T$ commutes with each $T_{\lambda_{i}}$. Thus
 $\ell$ $\leq\|T_{\lambda_{n}}\cdots T_{\lambda_{0}}g_{0}\|$ (by 15)
@@
 […]
-We conclude with some remarks concering Theorem $2.$
+We conclude with some remarks concerning Theorem 2.3.
 Remark 2.4 (without linearity, strong convergence fails).
-Thanks to an example constructed by Genel and Lindenstrauss (see [7]), it is known that the sequence $(x_{n})_{n\in{\mathbb{N}}}$ generated by the Krasnoselskii-Mann iteration may fail to converge strongly even when $\lambda_{n}\equiv\tfrac{1}{2}$.
+Thanks to an example constructed by Genel and Lindenstrauss (see [8]), it is known that the sequence $(x_{n})_{n\in{\mathbb{N}}}$ generated by the Krasnosel’skiĭ-Mann iteration may fail to converge strongly even when $\lambda_{n}\equiv\tfrac{1}{2}$.
 A posteriori, the underlying nonexpansive mapping cannot be linear.
@@
 Assume that the parameter sequence $(\lambda_{n})_{n\in{\mathbb{N}}}$ lies in $\left[0,\tfrac{1}{2}\right[$.
-Then the sequence $(x_{n})_{n\in{\mathbb{N}}}$ generated by the Krasnoselskii-Mann iteration satisfies
+Using mathematical induction, one checks that the sequence $(x_{n})_{n\in{\mathbb{N}}}$ generated by the Krasnosel’skiĭ-Mann iteration satisfies
 $x_{n}=\Big(\prod_{k=0}^{n-1}(1-2\lambda_{k})\Big)x_{0};$ (17)
-moreover, $x_{n}\to 0$ $\Leftrightarrow$ $\sum_{n\in{\mathbb{N}}}\lambda_{k}=\infty$.
+moreover, $x_{n}\to 0$ $\Leftrightarrow$ $\sum_{n\in{\mathbb{N}}}\lambda_{n}=\infty$.
 Therefore, if $\sum_{n\in{\mathbb{N}}}(1-\lambda_{n})\lambda_{n}<\infty$, then $(x_{n})_{n\in{\mathbb{N}}}$ does not converge to a point in $F$.
@@
 If it is a priori known that $T$ is averaged, then one may enlarge the range of parameter sequence to go beyond 11. (See [3, Section 3.1] for how this is done.)
-Remark 2.7 (an open problem).
-The original paper by Baillon, Bruck, and Reich [1] was situated in a more general setting with respect to the underlying space and the nonexpansive map; in fact, $T$ was merely assumed to be odd rather than linear.
-At present, we do not know if it is possible to extend Theorem 2.2 from linear to odd nonexpansive mappings.
-Acknowledgment
+3 Conclusions
+A well-known result by Reich on Krasnosel’skiĭ-Mann iterations yields weak convergence to a fixed point of a nonexpansive mapping $T$ under a general (divergent-series) assumption on the parameters.
+On the other hand, when $T$ is linear, work by Baillon, Bruck and Reich implies that the convergence is strong but the relaxation parameters needed to be constant.
+In this paper, we extend the Baillon-Bruck-Reich result from constant to general parameters in the linear case.
+The limit is explicitly identified, and some limiting examples and extensions are discussed.
+It would be interesting to know if this result remains true if $T$ is only assumed to be odd instead of linear as was the case in [1].
+Acknowledgments
+The authors thank the reviewers and the editor for constructive comments which significantly improved the manuscript.
 The research of SB was partially supported by a collaboration grant for mathematicians of the Simons Foundation.